OpenStudy (anonymous):

find all local max/min for f(x)=x^3-10x^2+20

OpenStudy (anonymous):

First find the derivative of f(x).. can you do that??

OpenStudy (anonymous):

wo3x-20x

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

Is it \(3x^2 - 20x\) ???

OpenStudy (anonymous):

lol my bad type error

OpenStudy (anonymous):

Now put this result to 0.. \[3x^2 - 20x = 0 \implies x(3x - 20) = 0 \implies x = 0 \qquad \And \qquad x = \frac{20}{3}\]

OpenStudy (anonymous):

So the two points we get are: x = 0 and x = 6.6 Draw it on number line: